# Observing the quantum Mpemba effect in quantum simulations

### v1

@@ II.3 Relaxation of tilted ferromagnetic states under dephasing
The experimental data are presented in Fig. 3 and the results are corroborated by numerical simulations. The measurement protocol is similar to the one presented in Fig. 1a except that the qubits are time-evolved under the ambient experimental noise. In this case, we observe the restoration of symmetry, however, the crossing of the EA for various tilt angles, unlike in the previous cases, is not observed. The experimental data and numerical simulations, see inset, imply that pure dephasing evolution does not show the QMPE while undergoing symmetry restoration of initially asymmetric states. In Fig. 3, solid lines correspond to numerical simulations, which are carried out for the experimentally measured dephasing rates listed in Methods IV.5.
II.4 Frobenius distance to observe QMPE
@@
Figure 4: Frobenius distances via classical shadows: the Frobenius distance (discussed in Methods IV.4) between the experimental time-evolved state and the theoretical equilibrium state is shown for three tilt angles and $N_{A}=4$. The state that is initially farthest from its equilibrium state attains the latter at the earliest time; confirming the QMPE in quantum simulations via Frobenius distance measures. Here also the solid lines are numerical simulations and the symbols are experimental data averaged over subsystems from the central region. Error bars are smaller than symbols.
@@
In a complementary approach, the route to symmetrization can also be probed by measuring the distance between two states — the experimental time-evolved state $\rho_{A}(t)$ and the diagonal ensemble $\rho_{A}^{\rm DE}$ [40] (more details in Methods IV.4), which describes the averaged behavior of any observable at long times. The diagonal ensemble is a mixed state that is diagonal in the eigenbasis of the quenching XY Hamiltonian and, therefore, it is $U(1)$ symmetric. According to the eigenstate thermalization hypothesis, for large systems, this ensemble is equivalent to a Gibbs ensemble [18, 19, 20]. From the data obtained in the randomized measurements, we evaluate the Frobenius distance [30] between $\rho_{A}(t)$ and $\rho_{A}^{\rm DE}$, see Methods  IV.4 for definitions.
@@
The results are presented in Fig. 4 for three tilt angles $\theta=0.2\pi,0.33\pi$ and $0.5\pi$. From the present analysis, it is clear that the states which start farther away from the corresponding diagonal ensemble attain it earlier than those which start comparably closer. As we see, the state at an angle $0.5\pi$ begins at the largest distance and relaxes to the diagonal ensemble the fastest. Such observation is the manifestation of the QMPE via Frobenius distance measurements.
Our findings constitute the first experimental evidence that the subsystem as a whole attains a stationary Gibbs ensemble. Although there exist previous experimental studies on the thermalization of isolated quantum systems [41, 42, 43, 44, 45, 46], they predominantly focused on specific local observables.
III Summary and outlook
@@
In this manuscript, we have presented the first experimental demonstration of the QMPE. Notably, this phenomenon manifests itself at times much shorter than those relative to finite-size effects, such as revivals. This separation of time scales enables us to observe the QMPE distinctly in our experimental setup, comprising of a modest number of qubits ($N=12$), even in presence of decoherence and disorder. Our investigation employs two distinct quantities, the entanglement asymmetry and the Frobenius distance, establishing that the QMPE is not specific to a particular observable. This versatility paves the way for further investigation employing a variety of theoretical and experimental tools.
@@
Our focus in this study has been on a specific class of initial states— the tilted ferromagnets. However, our findings motivate future studies with other configurations to determine the experimental conditions for the occurrence of the QMPE in ergodic systems. While the role of entanglement in driving thermalization in isolated quantum systems is well understood, it will be important to elucidate the interplay between entanglement and quantum fluctuations for the observation of the QMPE in generic quantum systems. The measurement protocol presented in this work is easily applicable to other platforms with individual control and readout capabilities such as arrays of atoms in optical lattices, Rydberg systems, or superconducting qubits [47, 48, 49, 17, 50]. This spotlights tantalizing opportunities for the experimental investigation of the QMPE, especially in higher dimensional systems where a richer phenomenology is expected due to the possibility of spontaneous symmetry breaking at finite temperatures. Our experimental findings may stimulate further theoretical investigations about the role played by dissipation and disorder in the dynamical restoration of symmetry. We believe that further studies on the QMPE will provide new protocols for faster preparation of thermal states (albeit at the subsystem level). Such states can then be used as input states for quantum simulation experiments. The experimentally accessible quantities discussed in the present manuscript, such as the EA and the Frobenius distance, will be valuable tools to assess the quality of the state preparation.
Note added:- While preparing our manuscript, we became aware of another experimental study of the Mpemba effect performed with a single trapped ion [51].
IV Methods
IV.1 Experimental set-up
@@
Here we use a trapped ion quantum simulator capable of storing and manipulating long ion strings for studying the QMPE. The experimental setup consists of a string of $N=12$ calcium ions that is made to emulate a spin-$1/2$ chain with long-range interactions. Each qubit is labeled by basis vectors $s=(\ket{\uparrow},\ket{\downarrow})\equiv(\ket{0},\ket{1})$, which are encoded onto |$\ket{\text{S}_{1/2},m=+1/2}$ and |$\ket{\text{D}_{5/2},m=+5/2}$ electronic levels of a calcium ion. The power-law XY interaction expressed in Eq. (2) is engineered with the help of a global laser beam that interacts with ions’ electronic and motional degrees of freedom from the transverse direction of the ion chain [16, 15]. Furthermore, the disordered Hamiltonian displayed in Eq. (3) is generated by shining a set of site-selective light shift beams along with the global entangling beam [39].
In our set-up, for the $x$ and $y$ global rotations, we use an elliptically shaped laser beam coupling to all ions simultaneously and driving the resonant optical transition between spin states. A tightly focused laser beam is used to perform local $z$ rotations, by inducing an AC-Stark shift of variable strength on individual ions. Quantum projective measurements in the $z$ basis are obtained via site-resolved fluorescence detection.
IV.1.1 AC-Stark shift compensation during the spin-spin interaction
@@
The trapped-ion platforms employ an Ising interaction when a bichromatic laser beam off-resonantly excites the motional red and blue sidebands of an ion chain [15]. In the case of symmetric detuning $\Delta$ of the bichromatic beam from the motional sidebands, an AC-Stark shift is present due to laser coupling to the ion’s spectator levels. For example, in our setup, when the interaction is generated on the quadrupole transition of a calcium ion chain with a 729-nm laser beam, the off-resonant coupling to the dipole-allowed transitions gives an AC-Stark shift. We cancel this shift by adding a third frequency component to the laser beam, such that the light off-resonantly couples to the quadrupole transition and induces an AC-Stark shift that is of opposite sign but of equal strength compared to the dipole transitions.
@@
The Ising interaction is then transformed into an XY interaction via asymmetrically driving the sidebands by $\Delta\pm\delta_{c}$. In the limit of $\delta_{c}$ being smaller than $\Delta$ and much larger than the spin-spin coupling strength $J_{0}$, the interaction Hamiltonian is written as [52],
@@
$H=\sum_{i>j}\frac{J_{0}}{2|i-j|^{\alpha}}\left(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{j}\sigma^{y}_{i}\right)+\frac{\delta_{c}}{2\Delta}\sum_{i}B_{i}\sigma^{z}_{i}~{},$ (4)
@@
where an additional AC-Stark shift is induced that is proportional to $\delta_{c}$. This term is often irrelevant when the input states are initialized in the eigenbasis of the $\sigma_{z}$ operator. However, it is relevant when the input state is not the eigenstate of the $\sigma_{z}$ operator, which is currently the case for $\theta\neq 0$. In the above equation, $B_{i}$ depends upon the motional occupation number $\langle N_{\ell}\rangle$, site-dependent Rabi frequency $\Omega_{i}$ and mode dependent Lamb-Dicke parameter $\eta_{i,\ell}$; and it is expressed as
@@
IV.2 Measurements
@@
The protocol described above is repeated for $N_{U}$ different random unitaries and recording $N_{M}$ projective measurements for each random unitary. The experimental data thus obtained comprises of random unitaries $U^{(r)}$ and the respective recorded bitstring measurements $\mathbf{s}^{(r,m)}=\left(s_{1}^{(r,m)},\dots,s_{N}^{(r,m)}\right)$ with $r=1,\dots,N_{U}$ and $m=1,\dots,N_{M}$ that are stored in a classical device for post-processing. From the collected RM dataset, we can construct operators known as classical shadows of the prepared state $\rho$ in the post-processing stage. This operator restricted to a given subsystem of interest $A$ can be defined as [29, 28]
$\hat{\rho}^{(r,m)}_{A}=\bigotimes_{i\in A}3\,{U_{i}^{(r)}}^{\dagger}\outerproduct{s_{i}^{(r,m)}}{s_{i}^{(r,m)}}U_{i}^{(r)}-\mathbb{I}_{2}~{}.$ (6)
@@
IV.3 Estimation of EA
The shadows defined in Eq. (6) can be used to extract arbitrary multi-copy functionals of the form $f_{n}=\mathrm{Tr}(O_{A}^{(n)}\rho_{A}^{\otimes n})$ defined on the subsystem of interest using the U-statistics estimator [67]:
$\hat{f}_{n}=\frac{1}{n!}\binom{N_{U}}{n}^{-1}\sum_{r_{1}\neq\dots\neq r_{n}}\mathrm{Tr}\left(O_{A}^{(n)}\bigotimes_{j=1}^{n}\hat{\rho}_{A}^{(r_{j})}\right)~{},$ (7)
which is an unbiased estimator of $f_{n}$, i.e $\mathbb{E}[\hat{f}_{n}]=f_{n}$. We also define here the classical shadow $\hat{\rho}_{A}^{(r)}=\mathbb{E}_{N_{M}}[\hat{\rho}_{A}^{(r,m)}]$, constructed by averaging over $N_{M}$ measured bitstrings for an applied unitary $U^{(r)}$. Additionally, the classical shadow formalism allows one to estimate interesting symmetry-resolved quantum properties of the underlying quantum state as shown in the Refs. [60, 68, 69]. In this work, in order to obtain an estimator $\widehat{\rm EA}$ of Eq. (1) from the RM dataset, we construct unbiased estimators of each of the individual trace moment term in Eq. (1) following Refs. [28, 68]. This writes as,
$\widehat{\mathrm{EA}}=\log\frac{\sum_{r_{1}\neq r_{2}}\mathrm{Tr}\left(\hat{\rho}_{A}^{(r_{1})}\hat{\rho}_{A}^{(r_{2})}\right)}{\sum_{r_{1}\neq r_{2}}\mathrm{Tr}\left(\hat{\rho}_{A,Q}^{(r_{1})}\,\hat{\rho}_{A,Q}^{(r_{2})}\right)}~{},$ (8)
where $\hat{\rho}_{A,Q}^{(r)}$ is the shadow for the symmetrized counterpart of $\hat{\rho}_{A}^{(r)}$ given by $\hat{\rho}_{A,Q}^{(r)}=\sum_{q\in\mathbb{Z}}\Pi_{q}\hat{\rho}_{A}^{(r)}\Pi_{q}$ that satisfies $\mathbb{E}[\hat{\rho}^{(r)}_{A,Q}]=\rho_{A,Q}$. Rigorous performance analysis of statistical errors from a finite number of experimental runs $N_{U}\times N_{M}$ in the estimation of such trace moments have been studied in various prior works [29, 59, 60, 63].
In the main text, the experiment for quenching with XY Hamiltonian (Fig. 1b) is done with $N_{U}=500$ and $N_{M}=30$ projective measurements for each random unitary. The experiment in the presence of disorder (results in Fig. 2) and pure dephasing (results in Fig. 3) is done with $N_{U}=300$ and $N_{M}=30$ projective measurements for each random unitary.
IV.4 Frobenius distance
For generic large quantum systems evolving unitarily, the reduced density matrix describing the subssytem $A$ is expected to relax, at long times, to a Gibbs distribution [18, 19, 20]. The eigenstate thermalization hypothesis [18, 19] predicts that such thermal state is equivalent to the reduced density matrix of the diagonal ensemble [40, 20]. The diagonal ensemble is obtained by projecting the initial density matrix onto the eigenbasis {|$\{\ket{E_{k}}\}$ of the quenching Hamiltonian, that is
@@ IV.4 Frobenius distance
where $\rho(\theta,t=0)$ is the initial state. The partial trace of this state $\rho_{A}^{\mathrm{DE}}(\theta)=\mathrm{Tr}_{\bar{A}}(\rho^{\mathrm{DE}}(\theta))$ gives the averaged behavior of local operators at long times. For each tilt angle $\theta$, we are interested in the Frobenius distance between the post-quench state of subsystem $A$ at a given time and the corresponding diagonal ensemble i.e., the distance FD$[\rho_{A}(\theta,t),\rho_{A}^{\mathrm{DE}}(\theta)]$, defined as [30],
$\mathrm{FD}(\rho_{A},\rho_{A}^{\mathrm{DE}})=\sqrt{1-\frac{2\mathrm{Tr}[\rho_{A}\rho_{A}^{\mathrm{DE}}]}{\mathrm{Tr}[\rho_{A}^{2}+(\rho_{A}^{\mathrm{DE}})^{2}]}}~{},$ (10)
where we have dropped the angle $\theta$ and time $t$ for brevity. In the main text, we utilized this distance to analyze the thermalization of the reduced density matrix after the XY Hamiltonian quench (see Fig. 4).
@@
IV.5 Experimental noise
In the disorder studies, we use a spin-echo sequence to mitigate slow drifts in the experimental parameters. To do so, an echo pulse, i.e. a $\pi$ rotation about the $x$ axis in the middle of the time evolution is applied. The coherence time measured is $T_{\mathrm{coh}}\sim 27(2)$ ms for strong disorders and $T_{\mathrm{coh}}\sim 27(4)$ ms for weak disorders. In the case of pure dephasing, where the prepared state is left to evolve under dephasing noise without interactions, we observe slow detrimental noise components affect the quality of randomized measurements. To overcome this, we again use a spin-echo sequence. The measured coherence time is $T_{\mathrm{coh}}\sim 45(4)$ ms.
In the numerical simulation shown in the results, we have considered a global dephasing mechanism with experimentally measured rates. The corresponding jump operator is $\sqrt{\Gamma}\sum_{i}\sigma_{i}^{z}$, where $\Gamma=1/(2T_{\mathrm{coh}})$. The spontaneous decay rate and depolarization noise rate reported elsewhere [12] are negligible compared to these rates and hence they are not the leading error sources for the current studies. To counteract slow drifts of the experimental parameters, the system is calibrated every few minutes for changes in clock transition, laser pulse area fluctuation, and shift in the ion position.
IV.6 Finite size effects: $4$-qubit experiment
Figure 5: EA in a 4-qubit system: EA, averaged over all possible subsystems, is plotted for $N_{A}=2$ and three tilt angles. The subsystem restores the symmetry and QMPE is observed. However, a revival of the EA is seen due to the finite-size effects.

### v2

@@ Abstract
The experimental data are presented in Fig. 3 and the results are corroborated by numerical simulations. The measurement protocol is similar to the one presented in Fig. 1a except that the states are time-evolved under the ambient experimental noise. In this case, we observe the restoration of symmetry, however, the crossing of the EA for various tilt angles is not observed. The experimental data and numerical simulations, see inset, imply that pure dephasing evolution does not show the QMPE while undergoing symmetry restoration of initially asymmetric states. In Fig. 3, solid lines correspond to numerical simulations, which are carried out for the experimentally measured dephasing rates listed in section C of SM.
Frobenius distance to observe QMPE–
@@
Figure 4: The Frobenius distance between the experimental time-evolved state and the theoretical equilibrium state is shown for three tilt angles and subsystem size $N_{A}=4$. The state that is initially farthest from its equilibrium state attains the latter at the earliest time; confirming the occurrence of the QMPE in quantum simulations via Frobenius distance measures. Solid lines are numerical simulations and the symbols are experimental data averaged over subsystems from the central region. The error bars are calculated via the jackknife resampling method (see SM section H)
@@
In a complementary approach, the route to symmetrization can also be probed by measuring the distance between two states — the experimental time-evolved state $\rho_{A}(t)$ and the diagonal ensemble $\rho_{A}^{\rm DE}$ Rigol et al. (2007). The diagonal ensemble describes the average behavior of any observable at long times. It is a mixed state that is diagonal in the eigenbasis of the quenching XY Hamiltonian and, therefore, is $U(1)$ symmetric. According to the eigenstate thermalization hypothesis, for large systems, this ensemble is equivalent to a Gibbs ensemble Deutsch (1991); Srednicki (1994); Rigol et al. (2008). From the data obtained in the randomized measurements, we evaluate the Frobenius distance Fagotti and Essler (2013) between $\rho_{A}(t)$ and $\rho_{A}^{\rm DE}$ (see more details in section G of SM)
@@
The results are presented in Fig. 4 for three tilt angles $\theta=0.2\pi,0.33\pi$ and $0.5\pi$. From the present analysis, we note that the states which start farther away from the corresponding diagonal ensemble attain it earlier than those which start comparably closer. As we see, the state at an angle $0.5\pi$ begins at the largest distance and relaxes to the diagonal ensemble the fastest. Such observation is the manifestation of the QMPE via Frobenius distance measurements. Our findings constitute the first experimental evidence that the subsystem as a whole attains a stationary Gibbs ensemble. Although there exist previous experimental studies on the thermalization of isolated quantum systems Gring et al. (2012); Trotzky et al. (2012); Kaufman et al. (2016); Neill et al. (2016); Ueda (2020); Zhou et al. (2022), they predominantly focused on specific local observables.
Conclusions and outlook–
@@
In this Letter, we have presented the first experimental demonstration of the QMPE. Notably, this phenomenon manifests itself at times much shorter than those relative to finite-size effects, such as revivals. This separation of time scales enables us to observe the QMPE distinctly in our experimental setup, comprising of a modest number of qubits ($N=12$), even in the presence of decoherence and disorder. Our investigation employs two distinct quantities, the entanglement asymmetry and the Frobenius distance, establishing that the QMPE is not specific to a particular observable. This versatility paves the way for further investigation employing a variety of theoretical and experimental tools.
@@
Our focus in this study has been on a specific class of initial states— the tilted ferromagnets. However, our findings motivate future studies with other configurations to determine the experimental conditions for the occurrence of the QMPE in ergodic systems. While the role of entanglement in driving thermalization in isolated quantum systems is well understood, it will be important to elucidate the interplay between entanglement and quantum fluctuations for the observation of the QMPE in generic quantum systems. The measurement protocol used in this work is easily applicable to other platforms with individual control and readout capabilities such as arrays of atoms in optical lattices, Rydberg systems, or superconducting qubits Daley et al. (2022); Bloch et al. (2008); Evered et al. (2023); Monroe et al. (2021); Hoke et al. (2023). This spotlights tantalizing opportunities for the experimental investigation of the QMPE, especially in higher dimensional systems where a richer phenomenology is expected due to the possibility of spontaneous symmetry breaking at finite temperatures. Our experimental findings may stimulate further theoretical investigations about the role played by dissipation and disorder in the dynamical restoration of symmetry. We believe that further studies on the QMPE will provide new protocols for faster preparation of thermal states (albeit at the subsystem level). Such states can then be used as input states for quantum simulation experiments. The experimentally accessible quantities discussed in the present manuscript, such as the EA and the Frobenius distance, will be valuable tools to assess the quality of the state preparation.
Acknowledgments–
The experimental team (JF, FK, MKJ, CR, RB) would like to acknowledge funding from the Institut für Quanteninformation GmbH and European Union’s Horizon 2020 research and innovation programme under grant agreement No 101113690 (PASQuanS2.1). LKJ, BV and PZ acknowledge funding from the Austrian Science Foundation (FWF, P 32597 N). LKJ acknowledges financial support from the PNRR MUR project PE0000023- NQST, and HPC and supercomputer facilities of the University of Innsbruck, where most of the numerical simulations were carried out. PZ acknowledges the funding from the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (651440). PC and FA are supported by ERC under Consolidator Grant number 771536 (NEMO). SM acknowledges the support from the Caltech Institute for Quantum Information and Matter and the Walter Burke Institute for Theoretical Physics at Caltech. Work in Grenoble is funded by the French National Research Agency via the JCJC project QRand (ANR-20-CE47-0005), and via the research programs Plan France 2030 EPIQ (ANR-22-PETQ-0007), QUBITAF (ANR-22-PETQ-0004) and HQI (ANR-22-PNCQ-0002). A.R. acknowledges support by Laboratoire d’excellence LANEF in Grenoble (ANR-10-LABX-51-01) and from the Grenoble Nanoscience Foundation. In numerical simulations, we have used the quantum toolbox QuTiP Johansson et al. (2013).
Supplemental material
.1 Experimental apparatus
@@
In this work, we used a trapped-ion quantum simulator capable of storing and manipulating long ion strings for studying the QMPE. The experimental setup consists of a string of $N=12$ 40Ca+ ions that is made to emulate a spin-$1/2$ chain with long-range interactions. Each qubit is labeled by basis vectors $s=(\ket{\downarrow},\ket{\uparrow})\equiv(\ket{1},\ket{0})$, which are encoded onto |$\ket{\text{S}_{1/2},m=+1/2}$ and |$\ket{\text{D}_{5/2},m=+5/2}$ electronic levels of a calcium ion. The power-law XY interaction expressed in Eq. 2 (see main manuscript) is engineered with the help of a global laser beam that interacts with ions’ electronic and motional degrees of freedom from the transverse direction of the ion chain Jurcevic et al. (2014); Porras and Cirac (2004). Furthermore, the disordered Hamiltonian displayed in Eq. 3 (see main manuscript) is generated by shining a set of site-selective light shift beams along with the global entangling beam Maier et al. (2019). In the experimental setup, we have noticed that the time-evolved states experience dephasing while undergoing the time evolution. The dephasing rates have been measured for all cases presented in this letter. The coherence times, which are related to dephasing rates via $\Gamma=1/(2T_{\mathrm{coh}})$, are $T_{\mathrm{coh}}\sim 16$ ms for the XY interaction case, $T_{\mathrm{coh}}\sim 45$ ms for the pure dephasing cases, $T_{\mathrm{coh}}\sim 27(2)$ ms for the strong disorders and $T_{\mathrm{coh}}\sim 27(4)$ ms for the weak disorder cases.
.2 AC-Stark shift compensation during the spin-spin interaction
@@
The trapped-ion platforms employ an Ising interaction when a bichromatic laser beam off-resonantly excites the motional red and blue sidebands of an ion chain Porras and Cirac (2004). In the case of symmetric detuning $\Delta$ of the bichromatic beam from the motional sidebands, an AC-Stark shift is present due to the coupling of the laser to the ion’s spectator levels. For example, in our setup, when the interaction is generated on the quadrupole transition of a calcium ion chain with a 729-nm laser beam, the off-resonant coupling to the dipole-allowed transitions gives an AC-Stark shift. We cancel this shift by adding a third frequency component to the laser beam, such that the light off-resonantly couples to the quadrupole transition and induces an AC-Stark shift that is of opposite sign but of equal strength compared to the dipole transitions.
@@
The Ising interaction is then transformed into an XY interaction via asymmetrically driving the sidebands by $\Delta\pm\delta_{c}$. In the limit of $\delta_{c}$ being smaller than $\Delta$ and much larger than the spin-spin coupling strength $J_{0}$, the interaction Hamiltonian is written as Arrazola et al. (2016),
@@
$H=\sum_{i>j}\frac{J_{0}}{2|i-j|^{\alpha}}\left(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{j}\sigma^{y}_{i}\right)+\frac{\delta_{c}}{2\Delta}\sum_{i}B_{i}\sigma^{z}_{i}\leavevmode\nobreak\ ,$ (4)
@@
where an additional AC-Stark shift proportional to $\delta_{c}$ is induced. This term is often irrelevant when the input states are initialized in the eigenbasis of the $\sigma_{z}$ operator. However, it is relevant when the input state is not the eigenstate of the $\sigma_{z}$ operator, which is currently the case for $\theta\neq 0$. In the above equation, $B_{i}$ depends upon the motional occupation number $\langle N_{\ell}\rangle$, site-dependent Rabi frequency $\Omega_{i}$ and mode dependent Lamb-Dicke parameter $\eta_{i,\ell}$; and it is expressed as
@@
.3 Experimental noise
SM Figure 1: We numerically examine the effect of global dephasing on EA and on the observation of the QMPE. Dashed lines correspond to unitary dynamics and solid lines correspond to dephasing dynamics. With increasing dephasing rates; a) $\Gamma=1$ rad/s, b) $\Gamma=10$ rad/s, and c) $\Gamma=100$ rad/s, a reduction in EA is observed for all three tilt angles. Conversely, the QMPE feature i.e. crossing of EA curves for three tilt angles is present even for the highest dephasing rate.
In the disorder studies, we use a spin-echo sequence to mitigate slow drifts in the experimental parameters. To do so, an echo pulse, i.e. a $\pi$ rotation about the $x$ axis in the middle of the time evolution is applied. The coherence time measured is $T_{\mathrm{coh}}\sim 27(2)$ ms for strong disorders and $T_{\mathrm{coh}}\sim 27(4)$ ms for weak disorders. In the case of pure dephasing, where the prepared state is left to evolve under dephasing noise without interactions, we observe that slow detrimental noise components affect the quality of randomized measurements. To overcome this, we again use a spin-echo sequence. The measured coherence time is $T_{\mathrm{coh}}\sim 45(4)$ ms.
In the numerical simulations shown in the results, we have considered a global dephasing mechanism with the experimentally measured rates. The corresponding jump operator is $\sqrt{\Gamma}\sum_{i}\sigma_{i}^{z}$, where $\Gamma=1/(2T_{\mathrm{coh}})$. The spontaneous decay rate and depolarization noise rate reported elsewhere Joshi et al. (2022a) are negligible compared to these rates and hence they are not the leading error sources for the current studies. To counteract slow drifts of the experimental parameters, the system is calibrated every few minutes for changes in clock transition, laser pulse area fluctuation, and shift in the ion position.
The impact of decoherence in observing the QMPE is further studied via numerical simulations. The results are displayed in Fig. 1 while considering various values of global dephasing rates $\Gamma$. We notice that features like the crossing of EA curves and observation of the QMPE remain unaffected. However, dephasing decreases EA for all tilt angles. The revival of EA is noticeably damped down by dephasing (see the peak at around $t=7.5$ ms). Nonetheless, we can conclude that for qualitative purposes the QMPE is robust under global dephasing.
.4 Implementation of disorder in the experiment
SM Figure 2: Disorder strength $h_{i}$ versus the ion index for the weak and strong disorder realizations employed in the experiments. Each color corresponds to one of the 5 sets of different realizations of disorders considered.
For the disorder studies presented in the main manuscript, the XY interaction is combined with site-dependent transverse field terms. This disorder Hamiltonian reads as
$H=\sum_{i>j}\frac{J_{0}}{2|i-j|^{\alpha}}\left(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{j}\sigma^{y}_{i}\right)+\sum_{i}h_{i}\sigma^{z}_{i}\leavevmode\nobreak\ ,$ (6)
where $h_{i}$ indicates transverse disorder strength. In our experimental system, these disorders terms are realized by shining multiple laser beams generated by an acousto-optic deflector (AOD) on individual ions. The AOD deflects multiple laser beams designated to individual ions with the desired light intensity such that site-dependent AC-Stark shifts are induced. We sample $5$ sets of weak and strong disorders randomly chosen from a uniform distribution function and apply them to the ion string. In Fig. 2, we show the measured disorders for both the strong and weak disorder cases.
.5 Measurements
@@
The protocol described above is repeated for $N_{U}$ different random unitaries and recording $N_{M}$ projective measurements for each random unitary. The experimental data thus obtained comprises of random unitaries $U^{(r)}$ and the respective recorded bitstring measurements $\mathbf{s}^{(r,m)}=\left(s_{1}^{(r,m)},\dots,s_{N}^{(r,m)}\right)$ with $r=1,\dots,N_{U}$ and $m=1,\dots,N_{M}$ that are stored in a classical device for post-processing. From the collected RM dataset, we can construct operators known as classical shadows of the prepared state $\rho$ in the post-processing stage. This operator restricted to a given subsystem $A$ can be defined as Huang et al. (2020); Elben et al. (2023)
$\hat{\rho}^{(r,m)}_{A}=\bigotimes_{i\in A}3\,{U_{i}^{(r)}}^{\dagger}\outerproduct{s_{i}^{(r,m)}}{s_{i}^{(r,m)}}U_{i}^{(r)}-\mathbb{I}_{2}\leavevmode\nobreak\ .$ (7)
@@
.6 Estimation of the EA
The shadows defined in Eq. (7) can be used to extract arbitrary multi-copy functionals of the form $f_{n}=\mathrm{Tr}(O_{A}^{(n)}\rho_{A}^{\otimes n})$ defined on the subsystem $A$ using the U-statistics estimator Hoeffding (1948):
$\hat{f}_{n}=\frac{1}{n!}\binom{N_{U}}{n}^{-1}\sum_{r_{1}\neq\dots\neq r_{n}}\mathrm{Tr}\left(O_{A}^{(n)}\bigotimes_{j=1}^{n}\hat{\rho}_{A}^{(r_{j})}\right)\leavevmode\nobreak\ ,$ (8)
which is an unbiased estimator of $f_{n}$, i.e $\mathbb{E}[\hat{f}_{n}]=f_{n}$. We also define here the classical shadow $\hat{\rho}_{A}^{(r)}=\mathbb{E}_{N_{M}}[\hat{\rho}_{A}^{(r,m)}]$, constructed by averaging over $N_{M}$ measured bitstrings for an applied unitary $U^{(r)}$. Additionally, the classical shadow formalism allows one to estimate interesting symmetry-resolved quantum properties of the underlying quantum state as shown in the Refs. Neven et al. (2021); Vitale et al. (2022); Rath et al. (2023). In this work, in order to obtain an estimator $\widehat{\rm EA}$ of Eq. 1 (main manuscript) from the RM dataset, we construct unbiased estimators of each of the individual trace moment term in Eq. 1 (main manuscript) following Refs. Elben et al. (2023); Vitale et al. (2022). This is written as,
$\widehat{\Delta S_{A}}=\log\frac{\sum_{r_{1}\neq r_{2}}\mathrm{Tr}\left(\hat{\rho}_{A}^{(r_{1})}\hat{\rho}_{A}^{(r_{2})}\right)}{\sum_{r_{1}\neq r_{2}}\mathrm{Tr}\left(\hat{\rho}_{A,Q}^{(r_{1})}\,\hat{\rho}_{A,Q}^{(r_{2})}\right)}\leavevmode\nobreak\ ,$ (9)
where $\hat{\rho}_{A,Q}^{(r)}$ is the shadow for the symmetrized counterpart of $\hat{\rho}_{A}^{(r)}$ given by $\hat{\rho}_{A,Q}^{(r)}=\sum_{q\in\mathbb{Z}}\Pi_{q}\hat{\rho}_{A}^{(r)}\Pi_{q}$ that satisfies $\mathbb{E}[\hat{\rho}^{(r)}_{A,Q}]=\rho_{A,Q}$. Rigorous performance analysis of statistical errors from a finite number of experimental runs $N_{U}\times N_{M}$ in the estimation of such trace moments have been studied in various prior works Huang et al. (2020); Elben et al. (2020); Neven et al. (2021); Rath et al. (2021b).
.7 Frobenius distance
For generic large quantum systems evolving unitarily, the reduced density matrix describing the subsystem $A$ is expected to relax, at long times, to a Gibbs distribution Deutsch (1991); Srednicki (1994); Rigol et al. (2008). The eigenstate thermalization hypothesis Deutsch (1991); Srednicki (1994) predicts that such a thermal state is equivalent to the reduced density matrix of the diagonal ensemble Rigol et al. (2007, 2008). The diagonal ensemble is obtained by projecting the initial density matrix onto the eigenbasis {|$\{\ket{E_{k}}\}$ of the quenching Hamiltonian, that is
@@
where $\rho(\theta,t=0)$ is the initial state. The partial trace of this state $\rho_{A}^{\mathrm{DE}}(\theta)=\mathrm{Tr}_{\bar{A}}(\rho^{\mathrm{DE}}(\theta))$ gives the average behavior of local operators at long times. For each tilt angle $\theta$, we are interested in the Frobenius distance between the post-quench state of subsystem $A$ at a given time and the corresponding diagonal ensemble i.e., the distance FD$[\rho_{A}(\theta,t),\rho_{A}^{\mathrm{DE}}(\theta)]$, defined as Fagotti and Essler (2013),
$\mathrm{FD}(\rho_{A},\rho_{A}^{\mathrm{DE}})=\sqrt{1-\mathrm{min}\left(1,\frac{2\mathrm{Tr}[\rho_{A}\rho_{A}^{\mathrm{DE}}]}{\mathrm{Tr}[\rho_{A}^{2}+(\rho_{A}^{\mathrm{DE}})^{2}]}\right)},$ (11)
where we have dropped the angle $\theta$ and time $t$ for brevity. Here, we have put a bound on the second term to bypass the unphysical Fidelities. In the limit of infinitely many measurements, the second term should always be upper bound by 1. In the main text, we utilized this distance to analyze the thermalization of the reduced density matrix after the XY Hamiltonian quench (see Fig. 4).
@@
.8 Estimation of error bars
For randomized measurements, experiments are carried out for $N_{U}$ random unitaries and $N_{M}$ number of projective measurements for each random unitary. In the main text, the time evolution with the XY interaction (Fig. 1b) is done with $N_{U}=500$ and $N_{M}=30$. For the inset of Fig. 1b, we sample $N_{U}=200$ and $N_{M}=100$. The experiments in the presence of disorder (results in Fig. 2) and pure dephasing (results in Fig. 3) are done with $N_{U}=300$ and $N_{M}=30$. We used a total of 5 disorder sets for each $\theta$ in the disorder studies.
In the main text, for the disorder case, the error bars are estimated from the standard deviation of the mean over disorder realization. Elsewhere, the error bars are estimated via the Jackknife method applied on the random unitaries that are used to estimate the averaged EA Wu (1986); Elben et al. (2023). Wherever the subsystem average data are presented, we employ the Jackknife resampling method by dropping one random unitary at a time and subsequently estimating the variance of the subsystem averaged EA. The subsequent error bars of the mean are presented in the figures.
.9 Finite size effects: $4$-qubit experiment
SM Figure 3: The EA after a quench with the XY Hamiltonian on a system of $N=4$ qubits. The averaged EA over all possible subsystems of size $N_{A}=2$ is plotted for three tilt angles. Symbols are the experimental data and curves correspond to numerical simulations. The subsystem restores the symmetry and the QMPE is observed. However, a revival of the EA is seen due to the finite-size effects.